Factor completely.
2b^2 + 200

Answer by jayanthihemkos@yahoo.com(87)

2b^2+100
2(b^2+50)
if the expression is equated to zero, then the value of each term is equated to zero
2(b^2+50)=0
b^2+50= 0
b^2=-50
b= square root of -50
b =iroot 50
b = 7.1i